### v1

# Kinematic and dynamical origins of mean-$p_{T}$ fluctuations in heavy-ion collisions

Abstract
Event-by-event fluctuations of the mean transverse momentum (mean-$p_{T}$) provide a sensitive probe of collective dynamics beyond single-particle spectra and anisotropic flow. In this Letter, we present a systematic comparison of mean-$p_{T}$ fluctuation measurements with calculations based on a Bayesian-calibrated multistage hydrodynamic framework. The experimental definitions employed by the STAR and ALICE Collaborations are implemented explicitly and found to yield consistent results within controlled limits. We study the centrality and beam-energy dependence of the observable, its sensitivity to key soft-sector ingredients, and the impact of the kinematic $p_{T}$ acceptance. By introducing scaled-$p_{T}$ cuts, we demonstrate that a part of the apparent energy dependence arises from kinematic projection effects, while the remaining trends reflect genuine collective dynamics. Our results establish mean-$p_{T}$ fluctuations as a nontrivial and independent validation of calibrated hydrodynamic descriptions of the quark–gluon plasma.
keywords:
transverse-momentum fluctuations , event-by-event correlations , relativistic heavy-ion collisions , multistage hydrodynamic models , kinematic acceptance effects
journal: Physics Letters B
[…]
1 Introduction
Relativistic heavy-ion collisions produce strongly interacting matter that exhibits pronounced collective behavior, commonly described as the hydrodynamic expansion of a quark–gluon plasma (QGP) Ollitrault:1992bk , Heinz:2013th . Extensive experimental and theoretical studies have shown that single-particle spectra and anisotropic flow coefficients provide stringent constraints on the bulk properties and transport coefficients of the medium Voloshin:2008dg , Song:2010mg , Bernhard:2019bmu , Nijs:2020ors , Nijs:2020roc , JETSCAPE:2020shq , JETSCAPE:2020mzn . Event-by-event fluctuations offer complementary information by probing the response of the system to initial-state inhomogeneities and dynamical fluctuations that are not accessible through event-averaged observables Jeon:2003gk , Gavin:2006xd .
Fluctuations of the event-wise mean transverse momentum are among the most interesting event-by-event observables. Experimental measurements have revealed characteristic centrality and beam-energy systematics, including an overall dilution of fluctuations with increasing system size and a nontrivial evolution from RHIC to LHC energies CERES:2003sap , STAR:2005vxr , STAR:2019dow , ALICE:2014gvd , ALICE:2024apz , ATLAS:2024jvf . These trends have been discussed in terms of partial thermalization, the number of effective particle-emitting sources, and the growth of collective radial expansion Gavin:2003cb , Gavin:2006xd , Gavin:2008ev , Bozek:2016yoj . Despite this long history, a quantitative interpretation of mean-$p_{T}$ fluctuations within a modern, globally calibrated multistage hydrodynamic framework has remained limited.
Recent progress in Bayesian parameter estimation Paquet:2023rfd has enabled the systematic calibration of multistage hydrodynamic models to a broad set of soft-sector observables, including particle yields, transverse-momentum spectra, and anisotropic flow coefficients Bernhard:2019bmu , Nijs:2020ors , Nijs:2020roc , JETSCAPE:2020shq , JETSCAPE:2020mzn . Such calibrations provide a controlled baseline for confronting additional observables that were not included in the fitting procedure. Mean-$p_{T}$ fluctuations are particularly well suited for this purpose, as they are sensitive to event-by-event variations of radial flow, freeze-out conditions, and late-stage hadronic dynamics, while being only weakly constrained by the observables typically used in the calibration.
An additional challenge in interpreting existing measurements is that different experiments employ distinct definitions and weighting procedures for mean-$p_{T}$ fluctuation observables. In particular, the prescriptions used by the STAR and ALICE Collaborations differ in their treatment of particle pairs, multiplicity weighting, and kinematic acceptance STAR:2019dow , ALICE:2014gvd , ALICE:2024apz . A meaningful model–data comparison therefore requires an explicit implementation of the experimental definitions, rather than relying on a single idealized observable.
In this Letter, we present a systematic comparison of mean-$p_{T}$ fluctuation measurements with predictions from a Bayesian-calibrated multistage hydrodynamic model at top RHIC and LHC energies JETSCAPE:2020shq , JETSCAPE:2020mzn . We implement the STAR and ALICE definitions explicitly, examine the centrality and beam-energy dependence of the observable, and study its sensitivity to key soft-sector model ingredients. We further investigate the role of the kinematic $p_{T}$ acceptance by comparing results obtained with experimental and scaled $p_{T}$ cuts, allowing us to disentangle genuine dynamical trends from kinematic projection effects.
The remainder of this Letter is organized as follows. Section 2 describes the hydrodynamic framework and the analysis setup, including the implementation of the experimental definitions of the transverse-momentum fluctuation observable. Section 3 presents the main results, focusing on the centrality and beam-energy dependence of mean-$p_{T}$ fluctuations, their sensitivity to key model ingredients, and the role of kinematic acceptance effects. Finally, Section 4 summarizes the main findings and outlines implications for future fluctuation measurements.
2 Model and analysis setup
2.1 Bayesian-calibrated multistage hydrodynamic framework
The model calculations presented in this work are based on the same Bayesian-calibrated multistage hydrodynamic framework JETSCAPE:2020shq , JETSCAPE:2020mzn , JETSCAPE:2022cob employed in our recent study of transverse-momentum–differential radial flow fluctuations, $v_{0}(p_{T})$ Du:2025dpu . Here we summarize only the essential ingredients relevant for the present analysis of mean-$p_{T}$ fluctuations and refer the reader to Refs. JETSCAPE:2020shq , JETSCAPE:2020mzn , JETSCAPE:2022cob , Du:2025dpu for a comprehensive description and validation.
The simulations are performed within the JETSCAPE framework (v3.7) Putschke:2019yrg using a TRENTo-based initial condition model Moreland:2014oya , followed by a short period of pre-equilibrium free streaming Broniowski:2008qk , Liu:2015nwa and subsequent second-order viscous hydrodynamic evolution with MUSIC Schenke:2010nt , Schenke:2011bn , Paquet:2015lta . The hydrodynamic stage includes temperature-dependent shear and bulk viscosities constrained by Bayesian inference to a broad set of soft-sector observables at both RHIC and LHC energies JETSCAPE:2020shq , JETSCAPE:2020mzn , JETSCAPE:2022cob . Hadronization is implemented via Cooper–Frye sampling at a switching temperature $T_{\mathrm{sw}}$ Cooper:1974mv , with viscous corrections to the distribution function Shen:2014vra , McNelis:2019auj , and the subsequent hadronic rescattering stage is modeled using the SMASH transport code SMASH:2016zqf .
In the present work, we use the Maximum A Posteriori (MAP) parameter set from the Bayesian analysis JETSCAPE:2020shq , JETSCAPE:2020mzn , JETSCAPE:2022cob as a fixed baseline and perform forward model calculations without retuning. This setup provides a predictive test of the calibrated framework at $\sqrt{s_{NN}}\!=\!200$ GeV and 5.02 TeV, since the mean-$p_{T}$ fluctuation observables at these energies were not included in the original calibration, while the corresponding data at $\sqrt{s_{NN}}\!=\!2.76$ TeV were JETSCAPE:2020shq , JETSCAPE:2020mzn . Variations around this baseline Du:2025dpu —such as switching off bulk viscosity, removing the hadronic afterburner, modifying the nucleon size in the initial conditions, or considering ideal hydrodynamic evolution—are used to assess the sensitivity of the observable to different stages of the collision dynamics.
2.2 Definition of the transverse-momentum fluctuation observable
The observable studied in this work is an event-by-event measure of correlated transverse-momentum fluctuations, designed to quantify dynamical fluctuations beyond trivial statistical sampling. Unlike single-particle spectra or average flow coefficients, this observable probes correlations between particles within the same event and is therefore sensitive to event-wise fluctuations of collective radial expansion, freeze-out conditions, and late-stage dynamics.
At the core of the analysis is the two-particle transverse-momentum covariance $C_{m}$, evaluated for particles within a given kinematic acceptance and centrality class. For a single event containing $N$ particles, the covariance is defined as Voloshin:2002ku , Voloshin:2003ud
$C_{m}^{(\mathrm{event})}=\frac{1}{N(N-1)}\sum_{i\neq j}(p_{T,i}-p_{T,\,\mathrm{ref}})(p_{T,j}-p_{T,\,\mathrm{ref}}),$ (1)
where $p_{T,i}$ denotes the transverse momentum of particle $i$ and $p_{T,\,\mathrm{ref}}$ is a reference mean transverse momentum. The precise definition of this reference mean distinguishes the experimental prescriptions used by different collaborations, as discussed below.
The experimentally reported quantity is obtained by averaging the event-wise covariance over all events within a given centrality bin, $C_{m}\!=\!\left\langle C_{m}^{(\mathrm{event})}\right\rangle_{\mathrm{events}}$, and normalizing it by the squared mean transverse momentum. Specifically, we define the dimensionless fluctuation measure
$R_{p_{T}}\equiv\sqrt{C_{m}}/\langle p_{T}\rangle,$
where $\langle p_{T}\rangle\!\equiv\!\langle[p_{T}]\rangle$ denotes the ensemble-averaged mean transverse momentum in the same centrality class and acceptance. This normalization removes trivial scaling with the absolute momentum scale and enables meaningful comparisons across centrality, beam energy, and particle species. [^1] Residual kinematic effects associated with the finite $p_{T}$ acceptance can still influence such comparisons, as discussed in Sec. 3.
Physically, $R_{p_{T}}$ quantifies the relative magnitude of correlated momentum fluctuations within an event. Event-by-event variations of the collective transverse velocity field lead to coherent shifts of particle momenta, generating positive two-particle covariances that survive the averaging over events. As a result, $R_{p_{T}}$ is particularly sensitive to fluctuations of radial collectivity and to mechanisms that amplify or damp such fluctuations, including viscous effects, initial-state granularity, and hadronic rescattering.
2.3 Implementation of experimental definitions
In this work, we explicitly implement the transverse-momentum fluctuation definitions used by the STAR and ALICE Collaborations at the particle level, applying identical kinematic cuts to the model and the data STAR:2019dow , ALICE:2014gvd , ALICE:2024apz . The two prescriptions differ in the choice of the reference mean $p_{T,\,\mathrm{ref}}$ appearing in Eq. (1). In the STAR definition, $p_{T,\,\mathrm{ref}}$ is taken to be the ensemble-averaged mean transverse momentum within a given centrality class, $p_{T,\,\mathrm{ref}}^{\mathrm{STAR}}=\langle[p_{T}]\rangle_{\mathrm{events}}$, where $p_{T}]$ is the event-wise mean transverse momentum. In contrast, the ALICE prescription uses the event-wise mean itself as the reference, $p_{T,\,\mathrm{ref}}^{\mathrm{ALICE}}=[p_{T}]^{(\mathrm{event})}$.
As shown analytically in A, the STAR observable can be decomposed into the ALICE-style two-particle covariance plus an additional term proportional to the variance of the event-wise mean transverse momentum, $\big\langle\big([p_{T}]_{i}-\langle p_{T}\rangle\big)^{2}\big\rangle$. The magnitude of this additional contribution depends on particle multiplicity and the width of the centrality bin, and may in principle lead to quantitative differences between the two definitions.
3 Results and discussion
3.1 Baseline comparison and sensitivity to soft kinematics
We begin with a baseline comparison of $\sqrt{C_{m}}/\langle p_{T}\rangle$ at $\sqrt{s_{NN}}=200$ GeV, shown in Fig. 1, where model calculations are confronted with STAR measurements as a function of centrality STAR:2019dow . The results are evaluated using the experimental kinematic acceptance $0.2<p_{T}<2$ GeV as in the data. Applying both the STAR and ALICE fluctuation definitions to the same set of model events with identical kinematic cuts, we find that the two prescriptions yield numerically indistinguishable results across all centralities considered. This confirms that, for the large multiplicities and narrow centrality bins employed here, the additional variance term discussed in A is subleading. In view of this near-equivalence, only the STAR-style result is shown in Fig. 1.
While the STAR- and ALICE-style model evaluations are mutually consistent, the model systematically underestimates the measured fluctuation magnitude when the experimental lower $p_{T}$ cut of $0.2$ GeV is used. A similar underestimation is observed in the Bayesian inference study at $\sqrt{s_{NN}}=2.76$ TeV, where mean-$p_{T}$ fluctuation data were included in the calibration, yet the MAP parameter set nevertheless yields systematically smaller fluctuation magnitudes for the experimental acceptance JETSCAPE:2020mzn . This indicates that the trend observed at 200 GeV reflects a systematic feature of the calibrated framework rather than an isolated discrepancy. To investigate the origin of this difference, we varied the lower edge of the acceptance window within the model while keeping the upper cut fixed. As illustrated in Fig. 1, lowering $p_{T}^{\min}$ leads to a noticeable increase of $\sqrt{C_{m}}/\langle p_{T}\rangle$ and improves the agreement with the data.
This sensitivity to the lower $p_{T}$ cut indicates that the $p_{T}$-integrated fluctuation observable is dominated by correlations carried by the very soft part of the spectrum, where collective radial dynamics are most pronounced. Schematically, the event-averaged covariance can be written as
$C_{m}\sim\int dp_{T}\,dp_{T}^{\prime}\,\frac{dN}{dp_{T}}\frac{dN}{dp_{T}^{\prime}}\,\langle\delta p_{T}\,\delta p_{T}^{\prime}\rangle,$ (2)
illustrating that the dominant contribution arises from the low-$p_{T}$ region where particle yields ($dN/dp_{T}$) are largest and momentum shifts induced by collective flow fluctuations are coherent (see also B). The residual discrepancy between the model and the data for the nominal experimental cuts therefore suggests that, within the present setup, the strength of event-by-event radial collectivity in the soft sector may be slightly underestimated. Crucially, the near-identical results obtained using the STAR and ALICE definitions demonstrate that this conclusion is robust with respect to reasonable variations in the fluctuation definition and is not an artifact of the averaging procedure.
These observations motivate the more systematic investigation of kinematic acceptance effects presented in the following sections, where we examine how the projection of momentum correlations onto different $p_{T}$ windows influences comparisons across beam energies and particle species.
Figure 1: Centrality dependence of $\sqrt{C_{m}}/\langle p_{T}\rangle$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. Model calculations are shown for two different lower transverse-momentum cuts, $0.15<p_{T}<2$ GeV and $0.2<p_{T}<2$ GeV, illustrating the sensitivity of the integrated observable to the kinematic acceptance. Published STAR data for $0.2<p_{T}<2$ GeV are shown for comparison STAR:2019dow .
3.2 Sensitivity to model ingredients at fixed beam energy
We next examine how $R_{p_{T}}\!\equiv\!\sqrt{C_{m}}/\langle p_{T}\rangle$ responds to variations of key model ingredients, focusing on Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. The results are summarized in Fig. 2, where the relative change of $R_{p_{T}}$ is shown with respect to a common baseline setup corresponding to the Bayesian-calibrated reference model JETSCAPE:2020shq , JETSCAPE:2020mzn , which includes shear and bulk viscosity, finite nucleon size (w=1.12w\!=\!1.12 fm), and a hadronic afterburner. To emphasize the intrinsic sensitivity of the observable, all results are presented as fractional deviations from this baseline.
Two representative centrality classes, 30–40% and 50–60%, are shown. This choice allows us to contrast a mid-central regime, where collective dynamics are well developed, with a more peripheral regime, where multiplicities are smaller and different fluctuation sources compete more strongly. As illustrated in Fig. 2, the response of $R_{p_{T}}$ to individual model modifications is clearly centrality dependent and does not follow a universal ordering across all variants.
Turning off bulk viscosity (“no bulk”) leads to a small change in $R_{p_{T}}$, with a slightly negative response in 30–40% collisions and a positive response in the more peripheral 50–60% class. This pattern suggests that bulk-viscous effects contribute more visibly to transverse-momentum fluctuations in dilute systems, where bulk-pressure variations are less efficiently damped. However, the overall magnitude of the response is modest in both centrality classes. Since turning off bulk viscosity hardens the single-particle spectra, it simultaneously enhances the correlated covariance sampled within the fixed $p_{T}$ window while increasing the mean transverse momentum in the denominator, leading to a partial cancellation in the normalized ratio $R_{p_{T}}$.
In contrast, the ideal-hydrodynamic limit (“ideal hydro”), in which shear viscosity is also switched off and viscous damping is absent, produces a systematic enhancement of $R_{p_{T}}$ in both centrality classes, with a larger relative effect in 50–60% collisions. This behavior is consistent with the interpretation of $R_{p_{T}}$ as a genuine fluctuation observable that is sensitive to dissipative smoothing of event-by-event radial expansion. A comparison of the “no bulk” and “ideal hydro” cases shows that $R_{p_{T}}$ responds more strongly to the removal of shear viscosity than to bulk viscosity within the present setup, indicating that shear-driven damping of transverse velocity gradients plays a dominant role in suppressing correlated momentum fluctuations.
This sensitivity pattern contrasts with that of the $p_{T}$-differential radial-flow fluctuation observable $v_{0}(p_{T})$, which has been found to be more sensitive to bulk viscosity and comparatively insensitive to shear effects Parida:2024ckk , Du:2025dpu , ALICE:2025iud , ATLAS:2025ztg . The difference reflects the distinct physical information encoded in the two observables. While $R_{p_{T}}$ integrates momentum correlations over a broad $p_{T}$ range and is therefore primarily governed by the overall coherence of event-by-event momentum shifts, making it particularly susceptible to shear-induced smoothing, $v_{0}(p_{T})$ resolves the momentum-dependent redistribution of radial-flow fluctuations and is thus more directly influenced by bulk-viscous effects that modify the spectral shape.
Additional insight is provided by removing the hadronic afterburner (“no afterburner”) which produces opposite trends in the two centrality bins: $R_{p_{T}}$ decreases in mid-central collisions but increases in peripheral ones. This behavior reflects a competition between correlation buildup during the hadronic stage and multiplicity-driven dilution effects, whose relative importance changes with system size.
Figure 2: Relative response of $R_{p_{T}}\!\equiv\!\sqrt{C_{m}}/\langle p_{T}\rangle$ to variations of model ingredients in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. Results are shown as fractional deviations with respect to the baseline calculation that includes shear and bulk viscosity, finite nucleon size, and a hadronic afterburner. Two representative centrality classes, 30–40% and 50–60%, are shown. The model variants correspond to switching off bulk viscosity (“no bulk”), disabling the hadronic afterburner (“no afterburner”), ideal hydrodynamic evolution without viscosity (“ideal hydro”), and a reduced nucleon size (“smaller nucleon”). Error bars indicate statistical uncertainties from the finite event sample.
Among the variations considered, reducing the nucleon size from w=1.12w\!=\!1.12 fm to $0.8$ fm (“smaller nucleon”) yields the largest and most robust enhancement of $R_{p_{T}}$ across both centralities, underscoring the strong sensitivity of the observable to initial-state granularity. A smaller nucleon width increases event-by-event fluctuations of local pressure gradients and hence amplifies fluctuations of the collective radial flow ($\beta_{T}$). At a schematic level, the correlated momentum shift induced by radial expansion can be expressed as $\delta p_{T}\sim m_{T}\,\delta\beta_{T}$, so that the two-particle covariance scales parametrically as
$C_{m}\sim\langle\delta p_{T}\,\delta p_{T}\rangle\sim\langle m_{T}^{2}\rangle\,\langle\delta\beta_{T}^{2}\rangle.$ (3)
Reducing $w$ enhances $\langle\delta\beta_{T}^{2}\rangle$ by increasing the granularity of the initial conditions, thereby strengthening collective momentum correlations and leading to a larger $R_{p_{T}}$.
Taken together, Fig. 2 demonstrates that $R_{p_{T}}$ is sensitive to multiple stages of the collision dynamics, including initial-state structure, viscous hydrodynamic evolution, and late-stage hadronic effects. The absence of a universal response pattern across centrality underscores that the observable probes competing fluctuation mechanisms whose relative contributions evolve with system size. Rather than a drawback, this feature makes $R_{p_{T}}$ a valuable complementary constraint on the modeling of event-by-event dynamics, particularly when combined with observables that primarily constrain average collective flow.
3.3 Impact of kinematic acceptance on the beam-energy dependence
In Fig. 1, we presented a baseline comparison between model calculations and STAR data at $\sqrt{s_{NN}}=200$ GeV, using the experimental kinematic window $0.2<p_{T}<2$ GeV. As a consistency check, we also examined the sensitivity of the result to variations of the lower $p_{T}$ cut within the model and found that modest changes of $p_{T}^{\min}$ lead to visible shifts in the magnitude of the integrated fluctuation observable. This observation already indicates that, for $p_{T}$-integrated correlation measures, the numerical value does not solely reflect the underlying dynamics, but also depends on how the momentum-space correlation structure is projected onto a finite kinematic acceptance. This projection effect must therefore be carefully controlled when comparing results across different beam energies.
This motivates a more systematic examination of kinematic effects when comparing results across different beam energies. In experimental measurements, the commonly used acceptance windows are not identical: at RHIC, STAR typically employs $0.2<p_{T}<2$ GeV at $\sqrt{s_{NN}}=200$ GeV STAR:2019dow , while at the LHC, ALICE uses $0.15<p_{T}<2$ GeV at $\sqrt{s_{NN}}=2.76$ and $5.02$ TeV ALICE:2014gvd , ALICE:2024apz . However, even if identical absolute $p_{T}$ cuts were applied at all energies, the corresponding momentum window would not probe the same relative portion of the underlying particle spectrum. As the characteristic transverse momentum scale of the system increases with beam energy, and varies with centrality at fixed energy, the same fixed $p_{T}$ range samples different regions of the soft momentum distribution and therefore different components of the underlying momentum correlations.
The physical origin of this effect can be understood from the structure of the single-particle transverse-momentum spectra themselves. For a fixed centrality class, we find that identified-particle $p_{T}$ spectra at different beam energies approximately collapse onto a common curve when expressed in terms of the dimensionless variable $p_{T}/\langle p_{T}\rangle$, up to an overall normalization factor. This approximate scaling indicates that the ensemble-averaged mean transverse momentum $\langle p_{T}\rangle$ provides a natural soft momentum scale of the system. Consequently, a fixed absolute $p_{T}$ window corresponds to different relative locations in scaled momentum space as the beam energy or centrality changes, even when the same nominal cuts are applied.
Figure 3: Beam-energy dependence of $\sqrt{C_{m}}/\langle p_{T}\rangle$ for charged hadrons in 30–40% central collisions, comparing results obtained with experimental kinematic cuts (solid symbols) and with scaled $p_{T}$ cuts (open symbols). The experimental cut setup follows the published experimental acceptances: $0.2<p_{T}<2$ GeV at $\sqrt{s_{NN}}=200$ GeV (STAR) and $0.15<p_{T}<2$ GeV at $\sqrt{s_{NN}}=2.76$ and $5.02$ TeV (ALICE). The scaled-cut results use energy-dependent windows defined relative to the charged-particle mean transverse momentum and anchored to the STAR acceptance at 200 GeV, yielding $0.226<p_{T}<2.264$ GeV at 2.76 TeV and $0.236<p_{T}<2.360$ GeV at 5.02 TeV.
Guided by this observation, we introduce a “scaled-cut” prescription designed to align the kinematic acceptance relative to a common soft momentum scale when comparing different beam energies. To disentangle genuine beam-energy systematics from kinematic projection effects, Fig. 3 compares two setups for mid-central (30–40%) collisions. The first corresponds to the “experimental cut” setup, in which the published acceptance windows are used at each beam energy. The second employs a scaled acceptance defined as
$p_{T}^{\min}(\sqrt{s_{NN}})$ $=\alpha\,\langle p_{T}\rangle_{\rm ch}(\sqrt{s_{NN}}),$ (4)
$p_{T}^{\max}(\sqrt{s_{NN}})$ $=\beta\,\langle p_{T}\rangle_{\rm ch}(\sqrt{s_{NN}}),$ (5)
where $\langle p_{T}\rangle_{\rm ch}$ is the charged-particle mean transverse momentum in the same centrality class. The constants $\alpha$ and $\beta$ are fixed by anchoring the scaled window to the STAR acceptance at $\sqrt{s_{NN}}=200$ GeV, $0.2<p_{T}<2$ GeV. This procedure yields energy-dependent momentum windows, $0.226<p_{T}<2.264$ GeV at 2.76 TeV and $0.236<p_{T}<2.360$ GeV at 5.02 TeV, that shift both the lower and upper edges upward at LHC energies, reflecting the harder underlying spectra.
Figure 3 shows that, under the experimental cuts, the integrated fluctuation measure exhibits an increase with beam energy. In contrast, when the scaled-cut prescription is applied, the results at both LHC energies are reduced relative to the fixed-cut case, with a larger suppression observed at $\sqrt{s_{NN}}=2.76$ TeV. This behavior naturally follows from the shift of the scaled acceptance toward higher transverse momenta, which samples a harder region of the momentum distribution than the published experimental cuts.
The reduction of $\sqrt{C_{m}}/\langle p_{T}\rangle$ under the scaled-cut prescription is closely connected to the $p_{T}$-cut dependence discussed in Sec. 3.1. There, lowering the minimum $p_{T}$ cut enhanced the fluctuation signal by including a larger fraction of very soft particles, where collective radial dynamics are strongest. In the present case, the scaled cuts act in the opposite direction at LHC energies by increasing both $p_{T}^{\min}$ and $p_{T}^{\max}$ relative to the experimental acceptance, thereby suppressing contributions from the softest, most strongly correlated part of the spectrum.
Within current uncertainties, the resulting beam-energy dependence under scaled cuts may even appear non-monotonic. We emphasize, however, that this behavior should not be over-interpreted as evidence for a genuine non-monotonic dynamical trend. Rather, it reflects a general kinematic projection mechanism inherent to $p_{T}$-integrated fluctuation observables. Because $R_{p_{T}}$ is dominated by correlations carried by the soft part of the spectrum, acceptance windows that shift toward higher relative momenta systematically suppress the measured fluctuation magnitude by reducing the weight of the most strongly correlated low-$p_{T}$ particles. Scaled cuts make this suppression explicit by construction, while fixed absolute $p_{T}$ cuts can induce it implicitly as the characteristic momentum scale of the system changes.
As a consequence, part of the observed beam-energy dependence of $p_{T}$-integrated fluctuation measures STAR:2019dow may arise from kinematic projection effects rather than changes in the underlying correlation strength itself (see B). When fixed absolute $p_{T}$ cuts are used across energies, the acceptance effectively drifts relative to the soft momentum scale of the system, altering the balance between soft and semi-soft contributions as the spectra harden or soften with $\sqrt{s_{NN}}$. The scaled-cut comparison presented here provides a controlled demonstration of this effect at top RHIC and LHC energies and highlights its direct relevance for RHIC Beam Energy Scan measurements STAR:2008med , STAR:2017sal , Bzdak:2019pkr , Du:2023efk , Du:2024wjm , where the progressive softening of the single-particle spectra can significantly enhance acceptance-induced biases in the apparent beam-energy systematics.
3.4 Species dependence and effective kinematic projection
Identified-particle measurements provide a more differential probe of transverse-momentum fluctuations by isolating their dependence on hadron mass and species composition. They offer a direct test of whether the observed correlations exhibit the mass dependence expected from fluctuating radial flow and help clarify how changes in particle composition may influence inclusive fluctuation measurements. At present, such species-resolved measurements of transverse-momentum fluctuations are not available experimentally, and the results presented here should therefore be viewed as exploratory predictions within the calibrated framework.
Figure 4 shows the fluctuation observable $\sqrt{C_{m}}/\langle p_{T}\rangle$ evaluated separately for identified $\pi^{\pm}$, $K^{\pm}$, and $p/\bar{p}$ in central (0–5%) Au+Au collisions at $\sqrt{s_{NN}}\!=\!200$ GeV, using a common kinematic window $0.2<p_{T}<2$ GeV for all species. For reference, the corresponding result for inclusive charged hadrons in the same centrality and acceptance is also indicated. A clear mass ordering is observed: the fluctuation measure increases systematically from pions to kaons to protons. Since the mean transverse momentum itself grows with particle mass, $\langle p_{T}\rangle_{p}>\langle p_{T}\rangle_{K}>\langle p_{T}\rangle_{\pi}$, this ordering implies an even stronger mass dependence of the absolute covariance $\sqrt{C_{m}}$. In other words, heavier particles not only have larger mean momenta, but also exhibit substantially stronger event-by-event momentum correlations.
This behavior is naturally understood as a consequence of fluctuating radial flow. Event-wise variations of the collective transverse velocity field lead to momentum shifts, $\delta p_{T}\sim m_{T}\,\delta\beta_{T}$, that scale approximately with particle mass, such that heavier hadrons experience larger absolute responses to the same underlying flow fluctuation. As a result, the two-particle covariance $C_{m}$, which encodes correlated momentum fluctuations, is amplified for heavier species. The observed mass ordering therefore provides further evidence that the dominant contribution to the fluctuation observable originates from collective dynamics rather than from few-particle or non-collective sources.
Figure 4: Species dependence of $\sqrt{C_{m}}/\langle p_{T}\rangle$ for identified $\pi^{\pm}$, $K^{\pm}$, and $p/\bar{p}$ in central (0–5%) Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV, evaluated in a common kinematic window $0.2<p_{T}<2$ GeV. The shaded band and dashed line indicate the corresponding result for inclusive charged hadrons in the same centrality and acceptance.
Importantly, the species dependence observed in Fig. 4 is closely connected to the kinematic acceptance effects discussed in Secs. 3.1 and 3.3. Although the same absolute $p_{T}$ window is applied to all species, this window corresponds to different regions of the underlying spectra when measured relative to each species’ characteristic momentum scale. Expressed in units of $p_{T}/\langle p_{T}\rangle$, the fixed window probes a softer and more flow-dominated region for protons than for pions, while sampling a comparatively harder region for lighter species. As a result, the fixed $p_{T}$ cut does not merely reveal the intrinsic mass dependence of the fluctuations but systematically amplifies it by enhancing the relative weight of soft, strongly correlated particles for heavier hadrons. Consequently, different particle species effectively project different parts of the underlying momentum-correlation structure even under identical kinematic cuts.
From this perspective, the mass ordering seen here represents an internal analogue of the explicit $p_{T}$-cut dependence discussed earlier. Varying the particle species at fixed acceptance shifts the effective location of the kinematic window in scaled momentum space in much the same way as changing the lower or upper $p_{T}$ cut for a single species. The enhanced sensitivity of heavier hadrons to both collective flow fluctuations and kinematic acceptance effects further reinforces the interpretation that the measured transverse-momentum correlations are dominated by global, event-wise dynamics of the expanding medium.
4 Conclusions
In this work, we have presented a systematic model study of event-by-event transverse-momentum fluctuations in relativistic heavy-ion collisions, focusing on the observable $R_{p_{T}}\equiv\sqrt{C_{m}}/\langle p_{T}\rangle$ and its sensitivity to soft-sector dynamics and kinematic acceptance. Using a Bayesian-calibrated event-by-event hydrodynamic framework, we performed baseline comparisons to RHIC data, investigated controlled variations of model ingredients, and examined the interplay between dynamical and kinematic effects in beam-energy systematics.
At $\sqrt{s_{NN}}=200$ GeV, the model reproduces the quantitative centrality dependence of the STAR measurements within the experimental acceptance. The remaining discrepancy in magnitude is traced to a pronounced sensitivity of $R_{p_{T}}$ to the very soft part of the transverse-momentum spectrum, indicating that the integrated fluctuation signal is dominated by long-wavelength collective dynamics carried by low-$p_{T}$ particles. This observation highlights the importance of kinematic acceptance in the interpretation of $p_{T}$-integrated fluctuation observables.
By varying individual model ingredients at fixed beam energy, we showed that $R_{p_{T}}$ responds to multiple stages of the collision evolution, including initial-state granularity, viscous hydrodynamic damping, and late-stage hadronic rescattering. The magnitude and even the sign of these responses depend on centrality, underscoring that transverse-momentum fluctuations probe competing fluctuation mechanisms whose relative importance evolves with system size. This behavior demonstrates that $R_{p_{T}}$ provides information complementary to average flow observables and offers an independent handle on event-by-event collective dynamics.
When comparing results across beam energies, we demonstrated that a substantial part of the apparent energy dependence of $R_{p_{T}}$ obtained under fixed experimental $p_{T}$ cuts can arise from kinematic projection effects. Introducing a scaled-cut prescription that aligns the acceptance relative to the characteristic soft momentum scale qualitatively modifies the observed energy dependence, emphasizing the need for acceptance-aware comparisons in RHIC-to-LHC studies and beam energy scan programs.
Finally, we explored the species dependence of transverse-momentum fluctuations in central Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. A clear mass ordering is observed, with heavier hadrons exhibiting stronger momentum correlations. This behavior is naturally explained by fluctuating radial flow and provides an internal analogue of the $p_{T}$-cut dependence discussed earlier, illustrating how different particle species effectively probe different regions of the underlying momentum-correlation structure even within identical kinematic windows.
Taken together, our results establish transverse-momentum fluctuation observables as sensitive probes of collective dynamics whose quantitative interpretation requires careful control of kinematic acceptance and projection effects. Future experimental measurements that explore alternative $p_{T}$ windows, identified-particle fluctuations, and differential extensions of the present analysis will provide valuable opportunities to further isolate dynamical contributions and to sharpen the sensitivity of $p_{T}$-correlation observables to the properties of the expanding quark–gluon plasma, in a manner complementary to traditional flow-based constraints.
Acknowledgements
The author acknowledges helpful discussions with workshop participants, especially I. Karpenko, R. Manikandhan and T. Reichert, at the workshop The QCD Critical Point: Are We There Yet? hosted by the Institute for Nuclear Theory (INT) at the University of Washington, which stimulated the present study. The author thanks the INT for its warm hospitality and stimulating research environment. This work was supported in part by the U.S. Department of Energy through the INT under Grant No. DE-FG02-00ER41132, and in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Grant No. DE-AC02-05CH11231. Computational resources were provided by the Ohio Supercomputer Center OhioSupercomputerCenter1987 . The author acknowledges the use of ChatGPT for assistance with grammar refinement and clarity improvement during manuscript preparation.
Appendix A Relation between the STAR and ALICE transverse-momentum fluctuation definitions
In Secs. 2.2 and 2.3, we introduced the transverse-momentum fluctuation observable $R_{p_{T}}=\sqrt{C_{m}}/\langle p_{T}\rangle$ and summarized the STAR and ALICE experimental prescriptions in a unified framework. In this appendix, we provide a compact analytical relation between the two definitions, clarifying why they yield nearly identical results in the present analysis.
For a given event $k$ with multiplicity $N_{k}$, we denote the event-wise mean transverse momentum by $p_{T}]_{k}=\frac{1}{N_{k}}\sum_{i}p_{T,i}$ and the ensemble-averaged mean by $\langle p_{T}\rangle=\langle[p_{T}]_{k}\rangle$. The STAR and ALICE definitions differ only in the choice of reference mean $p_{T,\mathrm{ref}}$ entering the event-wise covariance (Eq. (1)):
$p_{T,\mathrm{ref}}^{\mathrm{STAR}}=\langle p_{T}\rangle,\qquad p_{T,\mathrm{ref}}^{\mathrm{ALICE}}=[p_{T}]_{k}.$
Introducing the decomposition $p_{T,i}-\langle p_{T}\rangle=(p_{T,i}-[p_{T}]_{k})+([p_{T}]_{k}-\langle p_{T}\rangle)$, the STAR and ALICE single-event covariances, $C_{k}^{\mathrm{STAR}}$ and $C_{k}^{\mathrm{ALICE}}$, are related by the exact identity
$C_{k}^{\mathrm{STAR}}=C_{k}^{\mathrm{ALICE}}+\big([p_{T}]_{k}-\langle p_{T}\rangle\big)^{2}.$ (6)
Here $C_{k}$ denotes the event-wise covariance defined in Eq. (1), prior to the averaging over events that yields the experimentally reported quantity $C_{m}$. Averaging over events within a fixed centrality class yields
$\langle C_{k}^{\mathrm{STAR}}\rangle=\langle C_{k}^{\mathrm{ALICE}}\rangle+\mathrm{Var}([p_{T}]_{k}),$ (7)
where $\mathrm{Var}([p_{T}]_{k})\equiv\big\langle\big([p_{T}]_{k}-\langle p_{T}\rangle\big)^{2}\big\rangle$ is the event-by-event variance of the mean transverse momentum.
Equation (7) shows that the STAR definition contains, in addition to the ALICE-style two-particle covariance, an explicit contribution from fluctuations of the event-wise mean $p_{T}]$ within the centrality bin. As discussed in Sec. 2.3, the magnitude of this additional term depends on the particle multiplicity and the width of the centrality selection.
In the present analysis, the use of relatively narrow centrality bins and the large multiplicities characteristic of RHIC and LHC collisions render $\mathrm{Var}([p_{T}]_{k})$ numerically small compared to the genuine two-particle correlation contribution. Consequently, the STAR and ALICE prescriptions yield nearly identical values of $R_{p_{T}}$ when applied to the same set of model events with identical kinematic cuts, as explicitly demonstrated in Sec. 3.1. This near-equivalence justifies our use of a single representative definition in the presentation of results, while retaining both prescriptions as an internal consistency check on the robustness of the extracted fluctuation signal. However, in very peripheral collisions or for wide centrality selections, the variance term may become non-negligible.
Appendix B Kinematic projection effects and beam-energy systematics at RHIC BES
The interpretation of the beam-energy dependence of the transverse-momentum fluctuation observable $R_{p_{T}}=\sqrt{C_{m}}/\langle p_{T}\rangle$ requires careful separation of genuine dynamical effects from kinematic projection effects associated with the finite $p_{T}$ acceptance. This appendix provides a unified explanation of how fixed absolute $p_{T}$ cuts affect both the magnitude and the apparent energy dependence of $R_{p_{T}}$ at RHIC BES energies, and why a substantial part of the observed trend can arise from kinematics alone.
The two-particle covariance
$C_{m}\;\sim\;\int dp_{T}\,dp_{T}^{\prime}\left(\frac{dN}{dp_{T}}\right)\left(\frac{dN}{dp_{T}^{\prime}}\right)\langle\delta p_{T}\,\delta p_{T}^{\prime}\rangle$ (8)
is dominated by correlated momentum shifts generated by event-by-event fluctuations of collective radial expansion. To leading order, a fluctuation of the transverse flow velocity $\delta\beta_{T}$ induces a correlated shift of single-particle momenta, $\delta p_{T}\sim m_{T}\,\delta\beta_{T}$. Although higher-$p_{T}$ particles experience larger absolute momentum shifts, the rapidly falling single-particle yield, $dN/dp_{T}$, suppresses their statistical weight. As a result, the dominant contribution to $C_{m}$ arises from the low-$p_{T}$ region, where particle production is largest. This feature underlies the strong sensitivity of $C_{m}$ to the soft part of the spectrum.
At a fixed beam energy and centrality, modifying the $p_{T}$ acceptance therefore has a direct and predictable effect on $R_{p_{T}}$. Extending the acceptance toward lower $p_{T}$ includes additional particles from this correlation-dominated region, increasing the correlated covariance $\sqrt{C_{m}}$. At the same time, the ensemble-averaged mean transverse momentum $\langle p_{T}\rangle$ decreases due to the enhanced statistical weight of soft particles. Both effects act coherently to increase $R_{p_{T}}$. Conversely, restricting the acceptance to higher $p_{T}$ excludes soft correlated pairs, reduces $\sqrt{C_{m}}$, and increases $\langle p_{T}\rangle$, leading to a suppression of $R_{p_{T}}$. This behavior reflects a purely kinematic projection effect and does not imply any change in the underlying dynamics.
When a fixed absolute $p_{T}$ window is applied across different beam energies in the RHIC BES, an additional effect enters. As the beam energy is lowered, the characteristic momentum scale of the system decreases, and the same fixed $p_{T}$ window probes progressively harder regions of the spectrum when expressed in scaled units such as $p_{T}/\langle p_{T}\rangle$. This kinematic drift systematically removes the lowest-$p_{T}$ particles that carry the largest share of the correlated covariance. As a result, $\sqrt{C_{m}}$ decreases with decreasing beam energy even if the intrinsic strength of collective fluctuations were unchanged.
The denominator $\langle p_{T}\rangle$ is affected in the opposite direction: excluding soft particles increases the measured mean transverse momentum. This increase acts in the same direction as the suppression of $\sqrt{C_{m}}$, further reducing $R_{p_{T}}$; however, its quantitative impact is subleading compared to the loss of correlated soft pairs in the covariance. The correlated covariance $C_{m}$ is a two-particle quantity dominated by the multiplicity-rich soft sector, whereas $\langle p_{T}\rangle$ is a single-particle moment that varies more smoothly with acceptance. Consequently, the suppression of $\sqrt{C_{m}}$ dominates, and the net effect is a decrease of $R_{p_{T}}$ as the beam energy is lowered.
Across the wide energy range of the RHIC BES, this kinematic drift becomes pronounced. The fixed absolute $p_{T}$ cut moves rapidly toward the tail of the spectrum at lower energies, leading to a substantial loss of soft, collectively correlated particles. The resulting kinematic suppression of $R_{p_{T}}$ can therefore amplify the apparent beam-energy dependence observed in experimental measurements. Importantly, this mechanism operates independently of any genuine change in the microscopic fluctuation dynamics and must be accounted for when interpreting energy-dependent trends of $p_{T}$-integrated fluctuation observables.
Taken together, these considerations demonstrate that the decrease of $R_{p_{T}}$ toward lower RHIC BES energies obtained with fixed absolute $p_{T}$ cuts contains a significant kinematic component. A robust interpretation of beam-energy systematics therefore requires careful control of acceptance effects, either through scaled-cut prescriptions or through complementary differential measurements that explicitly resolve the momentum dependence of the underlying correlations.